OpenStudy (anonymous):

solve by substitution mthod y=2x-5 3x+y=15

OpenStudy (anonymous):

3x+(2x-5)=15 5x=20 x=4 so y =2x4-5=3

OpenStudy (anonymous):

3x + y = 15 3x + 2x - 5 = 15 5x = 20 x = 4 y = 3